### 03 Apr 2026 v2

# A Large-Scale Empirical Study of AI-Generated Code in Real-World Repositories

Abstract.
Large language models (LLMs) are increasingly used in software development, generating code that ranges from short snippets to substantial project components. As AI-generated code becomes more common in real-world repositories, it is important to understand how it differs from human-written code and how AI assistance may influence development practices. However, existing studies have largely relied on small-scale or controlled settings, leaving a limited understanding of AI-generated code in the wild.
In this work, we present a large-scale empirical study of AI-generated code collected from real-world repositories. We examine both code-level properties, including complexity, structural characteristics, and defect-related indicators, and commit-level characteristics, such as commit size, activity patterns, and post-commit evolution. To support this study, we develop a detection pipeline that combines heuristic filtering with LLM-based classification to identify AI-generated code and construct a large-scale dataset for analysis.
Our study provides a comprehensive view of the characteristics of AI-generated code in practice and highlights how AI-assisted development differs from conventional human-driven development. These findings contribute to a better understanding of the real-world impact of AI-assisted programming and offer an empirical basis for future research on AI-generated software.
1. Introduction
Recent advances in artificial intelligence, particularly large language models (LLMs), are profoundly reshaping modern software engineering practices. Systems such as OpenAI Codex, GPT-series models, and Claude, along with IDE-integrated assistants like GitHub Copilot and Cursor, have enabled developers to offload a wide spectrum of programming tasks to AI. These tools are now routinely used not only for generating small code snippets or completing functions, but increasingly for synthesizing entire files, modules, and even full-fledged projects. As a result, AI-assisted programming is transitioning from an experimental capability to a mainstream development paradigm, fundamentally altering how software is written and maintained.
This paradigm shift raises critical questions about the characteristics and implications of AI-generated code in real-world software systems. While early studies have examined LLM-generated code in controlled environments—typically focusing on small, self-contained programming tasks or benchmark datasets—such settings fail to capture the complexity and dynamics of real-world development. In practice, AI-generated code is often intertwined with human-written code, iteratively refined, and deployed within large, evolving codebases. Understanding the properties of such code therefore requires moving beyond lab-based evaluations toward comprehensive empirical studies in real-world repositories.
A key aspect of this understanding lies in analyzing code-level characteristics of AI-generated code. These include traditional software quality and complexity metrics, such as cyclomatic complexity, control-flow and data-flow complexity, loop nesting depth, and language-specific constructs (e.g., object-oriented metrics for Java or pointer-related complexity in C/C++). Additionally, defect-related indicators, such as bugs per KLOC, are essential for assessing reliability and maintainability. Comparing these metrics between AI-generated and human-written code can provide insights into whether AI assistance leads to simpler, more complex, or potentially more error-prone implementations.
Beyond code-level properties, it is equally important to investigate commit-level characteristics associated with AI-generated code. Modern software development is inherently collaborative and version-controlled, making commits a natural unit of analysis. Questions arise as to whether AI-assisted development changes developer behavior and workflow dynamics: Do AI-generated commits modify larger portions of code? Do they accelerate development by increasing commit frequency or reducing development time? How stable are such commits—do they require more frequent subsequent fixes, revisions, or reverts? Addressing these questions is crucial for understanding the broader impact of AI on software engineering processes.
Despite the growing interest in AI-assisted programming, there remains a significant research gap. Existing work primarily focuses on limited sets of code-level metrics and evaluates LLM-generated outputs in controlled, small-scale settings. To the best of our knowledge, no prior study has conducted a comprehensive, large-scale empirical analysis that jointly examines both code-level and commit-level characteristics of AI-generated code in real-world repositories. This gap limits our ability to draw reliable conclusions about the practical implications of adopting AI in software development.
In this paper, we aim to bridge this gap by conducting a comprehensive empirical study of AI-generated code in real-world projects. A central challenge in this endeavor is identifying and collecting AI-generated code from large-scale repositories, where such information is not explicitly labeled. To address this, we develop multi-stage rule-based message filter and LLM-assisted classifier for Github commit messages to construct a large dataset of AI-generated code instances. Based on this dataset, we perform, to the best of our knowledge, the most extensive measurement study to date, covering a wide range of code-level and commit-level metrics.
Our analysis yields a set of actionable insights into the current state of AI-assisted software development. These insights shed light on how AI-generated code differs from human-written code in terms of complexity, quality, and evolution, as well as how it influences development practices at the commit level. Such findings have important implications for researchers, practitioners, and tool designers seeking to better understand and improve the integration of AI into software engineering workflows.
In summary, this paper makes the following contributions:
$\bullet$ We build and release a large-scale (by far largest to our knowledge) dataset of AI-generated code extracted from real-world repositories, enabling various analytical study.
$\bullet$ Based on the dataset, we perform comprehensive measurement study of both code-level and commit-level characteristics, covering an extensive set of metrics, comparing AI-generated and human-written code.
$\bullet$ We distill multiple insights from our measurement findings regarding AI-assisted programming in practice and discuss potential implications for skateholders including programmars and AI vendors.
2. Related Works
Existing research on LLM-generated code can be broadly grouped into three lines: controlled prompt-based evaluation of generated outputs, empirical studies of real-world AI-assisted development, and detection or measurement of LLM-generated code in the wild.
Most prior work evaluates LLM-generated code in prompt-based or benchmark-style settings, where researchers design programming tasks, ask LLMs to generate code, and then analyze the resulting outputs. For example, (Molison et al., 2025) studied differences in maintainability and reliability between LLM-generated and human-written code, while (Li et al., 2026) examined performance and efficiency, and (Spiess et al., 2025) focused on call calibration and correctness of LLM-generated code. (Xu et al., 2026) compared code style differences between human-written and LLM-generated code and further tracked style trends in GitHub open-source projects. Other studies investigated code-level features such as line counts, cyclomatic complexity, nesting depth, and class counts (Rahman et al., 2025), while (Guo et al., 2026) examined lexical properties such as lexical density, however these features were often used mainly for classification or detection purposes, rather than for a systematic characterization of how LLM-generated code differs from human-written code across multiple structural dimensions.
Security is another major focus in prompt-based evaluations of LLM-generated code. Early work conducted exploratory assessments of the security of ChatGPT-generated code (Khoury et al., 2023). Subsequent studies expanded the evaluation to multiple models and further examined how model parameters and prompting strategies affect the security of generated code (Black et al., 2025). To enable more systematic assessment, (Siddiq et al., 2024) proposed a unified framework integrating code generation and security analysis, while (Mohsin et al., 2024) provided a more detailed study of the impact of prompt engineering on code security. (Sajadi et al., 2025) studied LLM-based security patching, which also sheds light on the security implications of LLM-generated code. In addition, (Wang et al., 2024) and (Yu et al., 2024) developed datasets and benchmarks for more systematic evaluation of the quality and security of LLM-generated code. Some work has further explored mitigation strategies, such as prompt engineering for safer code generation (Schaad et al., 2025) and model fine-tuning to improve the security of generated code (He and Vechev, 2023). Relatedly, several studies have investigated the use of LLMs themselves for vulnerability and bug discovery in code (Lin and Mohaisen, 2025; Chen et al., 2025). Overall, these studies provide important foundations for understanding the security of LLM-generated code, but they still predominantly evaluate model outputs under controlled settings.
Beyond controlled prompting, a growing body of work has begun to examine LLM-generated code in real-world open-source development. (Pearce et al., 2022) and (Fu et al., 2025) studied GitHub Copilot-generated code in real GitHub repositories, focusing on security issues and the ability of LLMs to repair vulnerabilities. (Xiao et al., 2026) analyzed self-admitted LLM usage in GitHub repositories, examining how generative AI is used in practice and surveying developers about usage policies. Using the AIDev dataset (Li et al., 2025) of LLM-agent pull requests, (Siddiq et al., 2026) investigated security-related issues, review interactions, and ecosystem characteristics, while (Hasan et al., 2026) examined cyclomatic complexity and code quality in the same setting. Other studies focused on specific aspects of AI-generated code in repositories, such as code smells in GitHub Copilot-generated code (Zhang et al., 2024) and adoption patterns reflected in modified lines and touched files at both commit and file levels for AI-agent-generated code (Robbes et al., 2026). These studies move beyond synthetic prompting scenarios and provide valuable evidence about how LLMs are actually used in practice.
Another related line of work studies the detection and measurement of LLM-generated code in the wild. (Daniotti et al., 2026) measured the usage of LLMs for coding by training a classifier for LLM-generated code, while (Suh et al., 2025), (Orel et al., 2025), and (Bisztray et al., 2026) also explored methods for detecting LLM-generated code. These studies are closely related because they help identify AI-generated code in large-scale repositories, but their main goal is detection or usage estimation rather than a systematic comparison of the characteristics of real-world LLM-generated code and human-written code.
Despite this growing body of work, an important gap remains. Prompt-based studies mainly evaluate generated outputs on designed tasks, which do not fully reflect how LLMs are used in real software development. Real-world repository studies, in contrast, usually focus on isolated aspects such as security issues, code smells, review interactions, or adoption patterns. As a result, there is still limited understanding of how real-world LLM-generated code systematically differs from human-written code, not only in code-level characteristics but also in development activity patterns.
Our work addresses this gap by conducting a systematic comparative study of real-world LLM-generated code and human-written code in open-source repositories. Rather than relying on crafted prompts or benchmark tasks, we analyze code that has actually been produced and integrated into real GitHub projects. Moreover, instead of focusing on a single aspect such as security, code smell, or adoption, we jointly examine both code-level characteristics and development-level characteristics, providing a broader empirical view of how LLM assistance affects real-world software development.
3. Measurement Pipeline
In this section we describe our measurement pipeline, from real-world dataset collection to the various metrics we consider and the methodology to measure them. The process of our pipeline is described in Figure 1.
Figure 1. Measurement Pipeline.
3.1. Dataset Collection and Alignment
As we target real-world code repositories, the first challenge is how to collect AI-generated code within them. To that end, we got inspiration from recent work (Schreiber and Tippe, 2025) that demonstrates the feasibility of identifying AI-generated code from GitHub by searching for explicit hints in code comments (e.g., “generated by ChatGPT”). However, such approaches suffer from several limitations. First, keyword search often introduces substantial noises due to substring matching (e.g., matching “use” in unrelated words), leading to false positives. Second, attribution comments are frequently ambiguous, as mentions of LLMs may refer to non-code tasks (e.g., documentation or data generation). Third, prior filtering strategies remain relatively coarse-grained and cannot reliably distinguish true AI-generated code from incidental references. To address these challenges, we design a refined data collection pipeline that improves both the precision and reliability:
Phase 1. Initial Data Collection. Following prior work (Schreiber and Tippe, 2025), we leverage the GitHub REST API to retrieve code files containing explicit mentions of popular LLM tools. We extend the scope to include recent models (e.g., Claude, Gemini, DeepSeek) and repositories up to 2025. After filtering for comments containing tool mentions, we obtain an initial dataset of 31,726 commits and 44,616 unique code files.
Phase 2. Multi-Stage Filtering. To improve data quality, we develop a multi-stage filtering pipeline:
(1) Keyword normalization. We eliminate false matches caused by substring-based search by ensuring that matched prefixes (e.g., “use”, “by”) appear as standalone words in comments. This step reduces noise introduced by the GitHub search mechanism.
(2) Semantic attribution filtering. We identify explicit attribution patterns (e.g., “generated by”, “written by”, “enhanced by”) and analyze their surrounding context to ensure that the LLM is used for code generation. Mentions referring to non-code artifacts (e.g., text, media, or data generation) are excluded.
(3) Usage-based exclusion. We remove cases where LLMs are mentioned only as tools for auxiliary tasks (e.g., API calls, analysis, or communication), rather than directly contributing to code generation.
After rule-based filtering, we obtain 4,882 commits and 8,141 files. Manual validation on randomly sampled data shows high precision (TP=88/100, TN=95/100), indicating the effectiveness of our rules.
Phase 3. LLM-assisted Classification. For samples that cannot be confidently classified by rules, we further employ an LLM-based classifier (ChatGPT 5-mini) to infer whether code is AI-generated. This step yields an additional 7,867 commits (10,713 files). Remaining uncertain cases are manually labeled, resulting in 686 additional positive samples. A secondary validation confirms the reliability of LLM-based classification (TP=85/100, TN=93/100).
Final Dataset. After deduplication using repository, file path, and commit hash, our final dataset contains 12,749 commits and 19,816 code files with confirmed AI involvement. To our knowledge, this is the largest dataset of real-world AI-generated code by far, which we will make public. Fig. 2(a) and Fig. 2(b) illustrate its distribution across LLM models/tools and programming languages, respectively. Consistent with prior observations (Schreiber and Tippe, 2025), general-purpose models (e.g., ChatGPT) dominate real-world usage, while code spans a diverse set of languages with Python and JavaScript being the most prevalent.
0%20%40%60%80%100%ChatGPT (83.3%)Copilot (10.8%)DeepSeek (2.2%)Gemini (1.7%)Cursor (1.3%)Others (0.7%)
(a) Distribution of AI-generated code by tools
0%20%40%60%80%100%Python (41.7%)JavaScript (15.4%)Java (8.6%)C++ (7.8%)TypeScript (7.8%)C# (4.8%)Swift (3.8%)C (3.7%)Others (6.4%)
(b) Distribution of AI-generated code by languages
Figure 2. Distribution of AI-generated code records by tools and programming languages.
Human Code Alignment. To enable fair comparison between AI-generated and human-written code, we construct a matched human code dataset. Direct comparison without control may introduce confounding factors, such as repository-specific coding practices, project complexity, or code scale differences. For example, variations in defect density or complexity metrics may stem from repository characteristics rather than the effect of AI assistance.
To mitigate such biases, we adopt an alignment strategy: for each AI-generated file, we randomly select two human-written files from the same repository and same programming language, with similar size (within 20% lines of code when possible). This design controls for repository-level and language-level factors, ensuring that observed differences more accurately reflect the impact of AI assistance rather than external confounders.
After these preprocessing, our final dataset contains 36,467 matched human written code files and 19,816 AI-generated files, enabling robust comparative analysis under controlled conditions.
3.2. Study Code-Level Characteristics
We aim at comprehensively understanding the differences between AI-generated and human-written code in real-world settings, to that end, we measaure an extensive and diverse set of code-level characteristics, surpassing prior works and providing a holistic comparison.
Metric Taxonomy. Specifically, our scope includes the following categories of metrics:
(1) Code Size and Basic Statistics. Basic properties such as lines of code (LOC), number of files, and token-level statistics, providing a baseline understanding of code scale and granularity.
(2) Structural Complexity. We analyze fine-grained structural components extracted from abstract syntax trees (ASTs), including functions, parameters, classes, statements, conditionals, loops, and recursion. We further derive complexity-related metrics such as cyclomatic complexity, nesting depth, and decision counts.
(3) Code Style and Duplication. Stylistic properties such as code duplication, reflecting maintainability and reuse patterns.
(4) Object-Oriented Design Metrics. For object-oriented languages (e.g., Java, C++, Python), we measure classical OOP metrics (e.g., class-level complexity and coupling) to understand design quality.
(5) Program Graphs. We analyze program characteristics from their control-flow, data-flow, and call graphs, capturing higher-level properties regarding both syntax and semantic.
(6) Code Defects and Security. We measure security-related properties such as vulnerability patterns and counts of potential defects, gaining insights of comparative code quality.
Metric Normalization. As the absolute volume of AI-generated and human-written code is different in our dataset (§3.1), direct metric comparison by raw counts is misleading. To address this, we normalize each metric using an appropriate measurement unit. For example, defect-related metrics are normalized by lines of code (e.g., bugs per KLOC), while structural metrics are measured at the function or file level (e.g., average number of loops per function).
Statistical Significance Testing. To draw meaningful and reliable conclusions in code characteristic differences betwene AI and human, we perform statistical hypothesis testing to determine whether observed AI/human differences are statistically significant.
Implementation. We employ various tools to compute the aforementioned metrics (e.g., cloc for basic statistics, CK and Understand for OOP metrics, CodeQL for security analysis, etc.). More details can be found in Table 1, which summarizes all measured metrics, along with their analysis tools and normalization scopes.
Table 1. Summary of code-level metrics, analysis tools, and measurement scopes.
Category Metrics Tool Scope
Code Size & Statistics LOC, token count, file size cloc (Danial, 2026), pygments (contributors, 2026) file / function
Structural Complexity #functions, parameters, statements, conditions, loops, recursion tree-sitter (tree-sitter contributors, 2025) per function / per file
Cyclomatic complexity, nesting depth, decision count AST-based analysis function
Code Style Code duplication ratio jscpd (jscpd contributors, 2026) file / project
OOP Design Class/method metrics (e.g., complexity, coupling) CK (Java) (Aniche, 2026), Understand (C++, Python) (SciTools, 2026) class
Program Graphs CFG/DFG properties, call graph size, dependencies joern (joern.io, 2026) + networkx (Hagberg et al., 2008) function / file
Security Vulnerability count, CWE patterns CodeQL (GitHub, 2025) per KLOC
Hard-coded secrets (passwords, API keys) String analysis file
3.3. Study Commit-Level Patterns
Beyond code-level characteristics (§3.2), it is equally important to understand how AI-generated code is produced, integrated, and evolved within real-world development workflows. Commit-level analysis provides a natural lens to study such dynamics, as commits capture both developer activity and project evolution. By examining commit histories associated with AI-generated code, we aim to uncover how developers collaborate with AI tools in practice and how AI-assisted code influences development processes and outcomes.
Research Questions. We specifically aim at answering the following research questions during our commit-level study, which collectively characterize the real-world usage patterns and lifecycle dynamics of AI-assisted code.
  • RQ1: Temporal patterns. Do AI-assisted commits exhibit different temporal characteristics (e.g., working hours, release proximity) compared to human-written code?
  • RQ2: Change characteristics. Do AI-generated commits involve different magnitudes or types of changes (e.g., number of modified files, lines of code, churn)?
  • RQ3: Post-commit evolution. Are AI-generated commits more likely to be revised, reverted, or modified after submission?
  • RQ4: Collaboration and workflow. How does AI-assisted code affect developer collaboration patterns, such as handoffs and multi-author contributions?
  • RQ5: Stability and maintenance. Does AI-generated code require more frequent fixes or exhibit different stabilization patterns compared to human-written code?
Methodology. To answer these questions, we extract commit-level information from real-world repositories corresponding to both AI-generated and human-written code. Specifically, information is collected from the following dimensions:
  • Temporal Characteristics (for RQ1). Commit timestamps, including time of day, day of week, and distance to the nearest release.
  • Change Characteristics (for RQ2). The scale and scope of code changes, including the number of modified files, lines, and overall churn.
  • Post-Commit Evolution (for RQ3, RQ4, and RQ5). The frequency of post-commit modification and revert within fixed time windows (e.g., 7, 30, and 90 days), the number of subsequent edits, participating authors, and time to first modification.
  • Workflow and Maintenance Patterns (for RQ4, and RQ5). These include fix likelihood, determined by whether subsequent commits contain fix-related keywords (e.g., “fix”, “revert”), stabilization time, measured as the time until the last modification of a file, and handoff latency, defined as the time until another developer modifies the code. We consider related issues in the interval of 90 days after the commit.
Statistical Hypothesis Testing. Similar to §3.2, we ensure that our comparative conclusions are statistically significant. Specifically, beyond hypothesis testing between AI-generated code and human written code, we also implemented the pairwise comparisons, that compare each AI-generated with one of the human written code in the same repository. we use Chi-squared tests (Pearson, 1900) for categorical variables and Mann-Whitney U tests (Mann and Whitney, 1947) for numerical variables in group-level analysis, and McNemar’s tests (McNemar, 1947) and Wilcoxon signed-rank tests (Wilcoxon, 1945) for paired comparisons.
Implementation. To collect the aforemntioned information and answer our research questions, we mainly use the git toolset with various corresponding commands (e.g., git diff for change characteristics and git log for some temporal metadata, etc.).
4. Code-Level Measurement Results
We present our measurement results and distilled findings regarding code-level metrics in this section, as outlined previously in §3.2.
4.1. Basic Code Statistics
Table 2 summarizes file-level basic statistics. Overall, AI-generated code exhibits consistently larger size across all LOC types. At the function level, token-based analysis (Table 3) reveals that AI-generated code uses more (unique) tokens, significantly more operators (Cohen’s $d=0.489$), operands, keywords, and identifiers. Furthermore, Table 4 shows that AI code also involves more per-function syntax constructs in general (except the switch-case constructs). These results suggest that rather than merely being larger, AI-generated code appears systematically more verbose across structural and non-structural elements, while the nearly identical comment ratio (Table 2) suggests that verbosity is not driven by documentation but by expanded code bodies.
Insight 1: Verbosity vs. Compression. AI-generated code is consistently more verbose across files and functions, not due to additional comments but due to expanded structural and token-level representations. In contrast, human-written code achieves similar functionality with more compact representations.
Table 2. Basic code statistics per file.
Metric AI Human
PLOC 256.57 192.68
SLOC 171.47 132.14
BLOC 38.87 25.93
CLOC 46.22 34.61
Blank ratio 15.15% 13.46%
Comment ratio 18.01% 17.96%
PLOC: physical lines; SLOC: source lines;
BLOC: blank lines; CLOC: comment lines
Another notable pattern is the contrast between content ratio and lexical density(Table 3). AI-generated code shows higher content ratio but lower lexical density, suggesting that it expresses more logic through repeated or explicit constructs rather than compact abstractions.
Insight 2: Explicitness over Compactness. AI-generated code favors explicit and expanded representations of logic (higher content ratio), while human-written code tends to compress semantics into denser expressions (higher lexical density).
Table 3. Code Token Statistics Distribution per Function
[…]
identifier count: amount of identifier tokens (e.g. variable, function, class names)
literal count: amount of literal tokens (e.g. numeric, string, boolean)
content ratio: (identifier count+literal count)/total token amounts
lexical density: ratio of unique tokens/total tokens
4.2. Structural Complexity
We next examine structural and control-flow complexity at both function and file levels.
Function-Level Structure.
Table 4 shows that AI-generated functions contain more statements (+2.61), conditions, loops, and slightly more parameters. However, effect sizes are generally small ($d<0.2$), suggesting consistent but moderate differences.
Insight 3: Expanded Local Structure. AI-generated functions tend to expand logic within functions by introducing more statements and control structures, leading to more explicit but not drastically more complex implementations.
Table 4. Code structure components statistics per Function
[…]
Control-Flow Complexity.
Table 5 shows that classical complexity metrics (cyclomatic complexity, decision points) are only slightly higher in AI-generated code, with very small effect sizes. However, a more meaningful difference emerges in nesting behavior. AI-generated code has higher overall and loop nesting depth, whereas human-written code exhibits slightly higher condition nesting and boolean complexity.
Insight 4: Layered vs. Compressed Complexity. AI-generated code tends to build complexity through deeper structural nesting (especially loops), whereas human-written code concentrates complexity within conditional logic and boolean expressions.
Table 5. Control Flow Metrics per Function
[…]
Halstead Complexity.
Table 6 reveals an interesting distributional pattern. AI-generated code has higher median vocabulary and length, indicating that typical AI functions are larger and more diverse. In contrast, human-written code shows significantly higher variance, with much larger maximum values.
Insight 5: Heavy-Tailed Human Complexity. AI-generated code is consistently moderate in size and complexity, whereas human-written code exhibits a heavy-tailed distribution with a small number of extremely large and complex functions.
Table 6. Halstead Metrics per Function
[…]
File-Level Structure.
Table 7 shows that human-written code contains substantially more functions per file (11.67 vs. 6.62), and these functions are also larger. In contrast, AI-generated code has slightly more methods and comparable class-level structure, suggesting a shift toward method-based organization.
Insight 6: Function-Oriented vs. Method-Oriented Design. Human-written code tends to organize logic into more and larger standalone functions, while AI-generated code favors method-based structuring within classes, indicating different decomposition strategies.
Additionally, recursive patterns are slightly less frequent in AI-generated code, suggesting a preference for iterative or explicitly expanded logic.
Insight 7: Reduced Use of Recursion. AI-generated code is less likely to use recursion, favoring iterative and explicitly expanded control flow.
Table 7. Code structure components statistics per File
Items AI codes Human codes
Function count 6.62 11.67
Function size(AST nodes) 138.97 201.82
Method count 4.95 4.59
Method size(AST nodes) 103.82 97.27
Class count 0.988 0.885
Class size(AST nodes) 584.91 574.59
Enumeration count 0.080 0.078
Recursive functions count 0.476 0.525
Recursive call count 0.587 0.672
4.3. Code Style
Code style reflects how developers organize, format, and reuse code. We examine two complementary aspects: duplication patterns and formatting structure.
Code Duplication
Table 8 summarizes duplication metrics. Overall, human-written code exhibits substantially higher duplication ratios than AI-generated code across all major measures (e.g., duplicated lines: 25.89% vs. 18.69%, duplicated intervals: 39.14% vs. 29.23%). This difference is primarily driven by cross-file duplication, where human-written code shows significantly higher reuse (24.52% vs. 17.20%). In contrast, within-file duplication remains low and similar across both groups.
Insight 8: Cross-File Reuse vs. One-Off Generation. Human-written code contains substantially more cross-file duplication, reflecting systematic reuse (e.g., shared utilities and templates), whereas AI-generated code is more often produced independently for each task, resulting in lower cross-file duplication.
Interestingly, AI-generated code exhibits a slightly higher number of clone instances per file (0.679 vs. 0.534), suggesting a different duplication pattern.
Insight 9: Fragmented vs. Structured Duplication. AI-generated code tends to produce more localized and fragmented duplication patterns, while human-written code exhibits fewer but larger and more structured duplicated regions.
Table 8. Code Duplication Results
Items AI codes Human codes
Duplicated lines ratio 18.69% 25.89%
Duplicated tokens ratio 19.71% 26.76%
Duplicated interval ratio 29.23% 39.14%
Clone instances(per file) 13455(0.679) 19477(0.534)
Within file duplicate lines ratio 1.86% 1.72%
Cross file duplicate lines ratio 17.20% 24.52%
Within file duplicate interval ratio 3.57% 3.23%
Cross file duplicate interval ratio 26.20% 36.54%
Formatting and Structural Style
Beyond duplication, formatting-related metrics reveal additional stylistic differences. As shown in Table 2, AI-generated code has a higher blank-line ratio (15.15% vs. 13.46%), while comment ratios are nearly identical.
Insight 10: Spacing-Oriented Formatting. AI-generated code uses slightly more blank lines, suggesting a tendency toward more spaced or segmented formatting, while human-written code is comparatively denser in layout.
The similar comment ratios (18.01% vs. 17.96%) indicate that AI-generated code does not rely more heavily on comments despite being larger in size.
Insight 11: Verbosity Without Additional Documentation. The increased size of AI-generated code is not driven by additional comments, but by expanded code structures, reinforcing the observation that AI favors explicit implementation over compact abstraction.
4.4. Object-Oriented Design Metrics
We analyze object-oriented design characteristics using standard CK-style metrics as listed in Table 9. Results are reported per class across C++, Java, and Python.
Language-dependent differences
A key observation is that differences between AI-generated and human-written code are strongly language-dependent. AI-generated object-oriented design is not uniformly simpler or more complex, but instead reflects different trade-offs across languages. The gap is largest in C++, minimal in Java, and moderate in Python, highlighting the importance of language-specific context when evaluating AI-generated code.
Insight 12: Language Sensitivity of OOP Differences. Differences in object-oriented design between AI-generated and human-written code vary significantly across languages, indicating that AI behavior is shaped by language-specific paradigms rather than a single consistent pattern.
Cross-language differences in object-oriented design
The object-oriented design differences between AI-generated and human-written code vary substantially across languages. In C++, human-written classes are much more complex and interconnected: WMC and CBO are both more than twice as large, and DIT is also notably higher, indicating deeper inheritance hierarchies and stronger inter-class dependencies. By contrast, Java shows near parity across almost all metrics. WMC, DIT, and RFC are nearly identical between the two groups, with only a slight reduction in CBO for AI-generated code, suggesting that AI-generated Java code largely follows established object-oriented design conventions. Python falls between these two cases: AI-generated classes exhibit slightly lower internal complexity (WMC, CBO) but slightly higher RFC, indicating somewhat simpler internal structures while preserving, or slightly expanding, externally reachable behavior.
Insight 13: Object-Oriented Differences Are Language-Dependent. The gap between AI-generated and human-written object-oriented design is not uniform across languages. It is most pronounced in C++, minimal in Java, and moderate in Python, suggesting that AI aligns more closely with mature and standardized design conventions, but diverges more in languages where deeper structural design and coupling patterns are common.
Table 9. OOP Metrics per Class
Items WMC DIT NOC CBO RFC LCOM
AI code C++ 15.79 0.209 0.226 1.22 9.76 41.34
Human code C++ 31.52 0.507 0.263 4.60 12.98 40.54
AI code Java 15.99 1.29 0.013 4.29 14.51 159.16
Human code Java 15.14 1.30 0.023 5.00 14.14 155.58
AI code Python 10.06 1.27 0.339 1.94 8.95 N/A
Human code Python 10.71 1.14 0.320 2.10 7.62 N/A
WMC: weighted methods per class; DIT: depth of inheritance tree;
NOC: number of children; CBO: coupling between objects;
RFC: response for class; LCOM: lack of cohesion.
4.5. Graph-Level Structural Analysis
We analyze program structure using graph-based representations, including control-flow graphs (CFG), data-flow graphs (DFG), and call graphs. These graphs capture how execution, data dependencies, and inter-procedural interactions are organized beyond surface-level syntax.
Overall observations
Across CFG and DFG metrics (Table 10 and Table 11), AI-generated code consistently exhibits slightly larger graphs (more nodes and edges) but lower density and centralization. These differences are statistically significant but small in magnitude.
Insight 14: Expanded but Less Centralized Graphs. AI-generated code tends to produce larger but less densely connected control-flow and data-flow graphs, suggesting a more distributed and stepwise representation of logic rather than tightly optimized structures.
This pattern aligns with earlier findings: AI-generated code expands logic into more explicit intermediate steps, resulting in additional nodes and edges while reducing structural concentration (e.g., lower PageRank and density).
Control-flow vs. data-flow consistency
The same structural trend appears in both CFG and DFG: AI-generated code has slightly higher node/edge counts and diameter, but lower density and centrality.
Insight 15: Consistent Expansion Across Abstraction Levels. The structural differences between AI and human code are consistent across both control-flow and data-flow representations, indicating that AI’s tendency toward explicit expansion is a general property rather than a metric-specific artifact.
Table 10. Control Flow Graph per Function
[…]
SCC: Strongly connected component; LWCC: largest weakly connected component
Table 11. Data Flow Graph per Function
[…]
Self Loop count $\pm0$ $\pm0$ / /
[…]
SCC and LWCC same with Table 10
Call graph structure
At the inter-procedural level (Table 12), the overall size of call graphs (nodes and edges) is comparable between AI and human code. However, AI-generated code shows lower density but greater depth and diameter.
Insight 16: Deeper but Not Broader Call Structures. AI-generated code does not introduce more functions or calls overall, but organizes them into deeper and more hierarchical call structures, suggesting a tendency toward layered decomposition rather than flat interaction patterns.
In contrast, human-written code shows slightly higher density and centrality, indicating more concentrated call interactions.
Interpretation
Taken together, these results reveal a consistent structural pattern: AI-generated code is not fundamentally more complex in terms of connectivity or centrality, but is more expanded and layered. Human-written code, in contrast, tends to concentrate logic into fewer, more interconnected structures.
Insight 17: Layered vs. Concentrated Structure. AI-generated code builds program structure through layered and distributed graph organization, while human-written code tends to concentrate complexity into more tightly connected regions.
Table 12. Call Graph per File
[…]
SCC and LWCC same with Table 10
4.6. Code Security and Defects
We analyze security and defect characteristics using static analysis (CodeQL) and manual inspection of sensitive information leakage. Metrics are normalized per k-lines to ensure fair comparison.
Overall security and defect density
As shown in Table 13, AI-generated code triggers slightly more total alerts than human-written code (12.81 vs. 11.58 per KLOC),which mainly driven by warnings and recommendations, whereas error rates are nearly identical between the two groups. From the risk perspective, AI-generated code contains fewer critical alerts, but noticeably more high-risk (0.934 vs. 0.464) and medium-risk alerts (0.282 vs. 0.182).
Insight 18: Higher Density of Security-Relevant Issues. AI-generated code does not show more critical or correctness-related failures overall, but it contains substantially more high- and medium-risk alerts. This suggests a denser concentration of security-relevant issues, even if the most severe category does not increase.
Table 13. CodeQL Alerts per k-Lines
Items AI codes Human codes
Total Alerts 12.81 11.58
Errors 0.840 0.846
Warnings 2.41 2.23
Recommendations 8.29 7.80
Critical Security Alerts 0.042 0.052
High Security Alerts 0.934 0.464
Medium Security Alerts 0.282 0.182
Vulnerability composition
Table 14 shows that both groups share the same dominant CWE categories (CWE-563, CWE-561, CWE-396), suggesting similar primary sources of issues. Differences emerge in secondary categories: AI-generated code shows relatively more input validation (CWE-20), resource consumption (CWE-400), and logging-related issues (CWE-117), while human-written code shows more concurrency, resource lifecycle, and logical condition issues (e.g., CWE-208, CWE-772, CWE-570).
Insight 19: Shifted Vulnerability Composition. AI-generated and human-written code share similar dominant defect types, but differ in secondary distributions, suggesting that AI shifts the composition of common issues rather than introducing entirely new vulnerability patterns.
Table 14. CodeQL high frequency CWEs
Rank AI code CWE ratio(count) Human code CWE ratio(count)
1 CWE-563 32.3%(8379) CWE-563 34.4%(9976)
2 CWE-561 9.08%(2358) CWE-561 13.4%(3904)
3 CWE-396 6.28%(1630) CWE-396 6.42%(1863)
4 CWE-20 2.88%(747) CWE-208 3.63%(1053)
5 CWE-390 2.86%(742) CWE-772 3.33%(968)
6 CWE-400 2.63%(683) CWE-390 3.00%(870)
7 CWE-772 2.38%(619) CWE-248 2.26%(656)
8 CWE-208 2.36%(614) CWE-570 2.16%(626)
9 CWE-117 2.34%(608) CWE-571 2.16%(626)
10 CWE-248 2.19%(568) CWE-20 2.06%(597)
Language-dependent security patterns
Table 15 reveals strong language-specific effects. AI-generated code shows higher alert density in C, C++, C#, and especially JavaScript (13.98 vs. 6.25), while human-written code is higher in Java, Python, Ruby, and TypeScript. The difference is particularly pronounced for high-risk alerts: JavaScript shows a large increase in AI-generated code (4.02 vs. 1.14), whereas TypeScript shows the opposite trend (0.078 vs. 0.794).
Insight 20: Language-Dependent Security Impact. The security impact of AI-generated code varies substantially across languages. In particular, JavaScript shows a clear increase in high-risk vulnerabilities, while TypeScript exhibits the opposite pattern, highlighting ecosystem-specific effects.
Table 15. CodeQL Alerts Language Distribution per k-lines
Language AI Alerts Human Alerts AI high risk Human high risk
C 4.95 4.28 0.904 0.420
C++ 5.64 3.88 0.053 0.049
C# 13.25 12.20 0 0
Java 15.12 16.32 0.357 0.238
Javascript 13.98 6.25 4.02 1.14
Python 18.63 21.71 0.551 0.453
Ruby 0.241 1.29 0.241 1.29
Typescript 4.92 5.68 0.078 0.794
Sensitive information leakage
Table 16 shows comparable overall leakage levels but different distributions. Human-written code contains more password-related leaks, while AI-generated code shows higher rates of API keys, tokens, and private keys.
Insight 21: Different Secret Leakage Patterns. Human-written code is more prone to password-like leakage, whereas AI-generated code more frequently exposes token- and key-based credentials, suggesting different leakage mechanisms and mitigation needs.
Table 16. Private Message in codes per 10k lines
Items AI code count Human code count
Hardcode Password 0.703 1.02
Simple Password 0.512 0.811
API key or token 0.194 0.098
Private key 0.021 0.006
Simple password: Pure number or letters, short password,
contains ’password’ itself or varieties
5. Commit-Level Measurement Results
Table 17. Results of Github Commit analysis
Items AI values Human values p-value q-value 1/2
Non-work Hours 0.372 0.358 0.001 0.000 / 0.000
Weekend Commit 0.232 0.237 0.190 0.753 / 1.000
Merge Commit 0.126 0.000 0.000 0.000 / 0.000
Release in 30 days 0.622 0.694 0.000 0.000 / 0.000
[…]
Files touched 7 days 0.004 0.002 0.001 0.243 / 1.000
Files touched 30 days 0.013 0.003 0.000 0.000 / 0.000
Files touched 90 days 0.020 0.003 0.000 0.000 / 0.000
[…]
Revert in 90 days 0.000 0.000 0.040 0.286 / 0.588
[…]
Table 17 summarizes commit-level comparisons between AI-associated and human-written code, and we report both pooled group-level and matched pair-wise results. The $p$-value compares all AI-associated samples with all human-control samples as two overall groups, while the $q$-value is derived from matched pair-wise comparisons between each AI-associated sample and its matched human-control commit; q-value 1 and q-value 2 denote the first and second matched controls, respectively. Throughout this paper, values below 0.05 are considered statistically significant. Therefore, significance in both $p$ and $q$ indicates a robust difference; significance only in $p$ indicates a group-level difference that weakens after matching; significance only in $q$ indicates a difference that is clearer after matching; and significance in neither indicates no reliable difference. Results are organized by the research questions introduced in §3.3.
RQ1: Temporal patterns.
Commits contains AI-associated contents are slightly more likely to occur during non-working hours (0.372 vs. 0.358), while weekend activity shows no consistent difference under paired comparisons. In addition, commits with AI-associated codes are less likely to occur within 30 days before a release (0.622 vs. 0.694).
Insight 22: Flexible but Risk-Aware Usage. Commits with AI-assisted codes are more likely to occur outside regular working hours, but are less frequent near release deadlines, suggesting that developers use AI flexibly while avoiding it in stability-critical stages.
RQ2: Change characteristics.
AI-associated commits are substantially smaller in scope. They involve fewer files and significantly lower line churn per commit compared to human-written code.
Insight 23: Localized and Fine-Grained Edits. AI-assisted coding is primarily used for small, localized modifications rather than large-scale, multi-file changes, indicating a role as a fine-grained development aid.
RQ3 & RQ5: Post-commit evolution and maintenance.
Differences are limited in the short term (7 days), but become pronounced over longer horizons. AI-generated code is more likely to be revisited within 30 and 90 days, with higher touch frequency, more participating authors, and greater accumulated churn. It also shows slightly higher likelihood of follow-up fixing commits and significantly longer stabilization time. However, most median values remain zero, indicating that these effects are driven by a subset of commits with repeated follow-up activity.
Insight 24: Delayed but Concentrated Maintenance. AI-generated code does not exhibit substantially higher immediate instability, but it incurs more follow-up modifications and longer stabilization over time, suggesting increased downstream adjustment costs concentrated in a subset of cases.
RQ4: Collaboration and workflow.
AI-associated commits involve longer handoff latency (28.28 vs. 11.23 days) and more follow-up contributors over time, indicating a different collaboration pattern. At the same time, merge commits appear more frequently in AI-associated changes.
Insight 25: Slower Handoffs and Distributed Follow-Up. AI-assisted code tends to remain longer with the original author before being handed off, but eventually involves more distributed follow-up activity, suggesting a shift toward delayed but broader collaboration.
6. Discussion
Threats to Validity. We discuss some potential threats to validity of our work as follows:
$\bullet$ Construct validity is challenged by the difficulty of identifying AI-generated code in the wild, as AI involvement is rarely explicitly labeled and there is no large-scale real-world dataset. We mitigate this using a multi-stage filtering pipeline as described in §3.1, though false positive and negatives can still occur regardlessly, our filtering achieves high accuracy in sampled verification and ensures high-fidelity statistical study.
$\bullet$ Internal validity may be affected by repository-level confounders. We reduce this risk by aligning AI and human samples (§3.1) within the same repository, language, and similar code size, though residual differences may remain.
$\bullet$ Measurement validity depends on the accuracy and coverage of analysis tools. While we employ well-established tools and normalized metrics, tool limitations and language-specific behaviors may introduce noise. Therefore, we emphasize consistent cross-metric trends rather than conclusions from any single metric.
$\bullet$ Statistical conclusion validity is influenced by the large sample size, which can make small effects statistically significant. We interpret results using both statistical significance and effect size, and complement group-level analysis with paired comparisons where applicable.
$\bullet$ External validity is limited to public GitHub repositories with explicit AI-related signals, and is dominated by certain tools and languages (e.g., ChatGPT, Python, JavaScript). While this may limit generalization, it reflects real-world open-source usage and provides an in-vivo perspective on AI-assisted development.
Scope and Interpretation of the Results. Unlike prior work that evaluates LLM-generated code under controlled prompts, we study the observable outcome of real-world AI-assisted programming: code that has been selected, edited, and integrated into real repositories. Our goal is therefore not to characterize all possible model outputs—which are inherently prompt-dependent—but to understand the typical properties of AI-involved code in practice.
Under this scope, a consistent pattern emerges: AI-involved code tends to be more explicit, more verbose, and more layered. It is generally larger, uses more structural constructs, and exhibits deeper nesting, while AI-associated commits are smaller, more localized, and more likely to incur follow-up modification over time. These findings reflect the combined effect of model behavior and human integration, rather than raw model outputs alone.
Implications for Developers and AI Vendors. Our results suggest that AI assistance reshapes development effort rather than simply reducing it. While AI-generated code is effective for localized edits and rapid drafting, it is associated with increased follow-up modification and longer stabilization over time. Developers should therefore anticipate a shift of effort from initial coding to downstream validation and maintenance, and allocate review and testing effort accordingly beyond the immediate post-commit window.
More importantly, our findings indicate that AI-generated code tends to be more explicit, less compact, and more layered. This suggests that the primary risk is not correctness failure, but structural inefficiency—including unnecessary verbosity, reduced reuse, and fragmented logic. Developers may benefit from treating AI outputs as drafts to be consolidated, focusing on abstraction, deduplication, and structural simplification after generation.
For AI tool vendors, these results highlight a gap between generation quality and long-term code quality. Current systems optimize for immediate correctness and completion, but less for maintainability. Our findings suggest that future tools should better support post-generation refinement, such as encouraging reuse across files, reducing unnecessary expansion, and warning about overly deep or fragmented structures. In addition, the strong language-dependent differences observed in our study indicate that AI assistants should adopt more language-aware optimization strategies rather than relying on uniform generation behavior.
Conclusion
In this paper, we presented a large-scale empirical study of AI-generated code in real-world repositories. Unlike prior work based on controlled prompts, we analyze code that has been selected, edited, and integrated into real projects, and compare it with matched human-written code. Our results show that AI-generated code is generally more verbose, more explicit, and more structurally layered, while human-written code is more compact and relies more on reuse. At the process level, AI-associated commits are smaller and more localized, yet incur more follow-up modification and longer-term maintenance. These findings suggest that AI assistance does not simply change who writes code, but reshapes how code is expressed and evolved. It acts as a local productivity aid while shifting effort toward downstream refinement. By releasing our measurement pipeline, we provide a foundation for future studies to track the evolving impact of AI on software engineering in practice.